40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 6 01:35:03 2012, kiwamu, Update, LSC, MICH characterization 
    Reply  Tue Mar 6 23:14:38 2012, rana, Update, SUS, Optical Lever had bad filters Untitled.png
Message ID: 6367     Entry time: Tue Mar 6 23:14:38 2012     In reply to: 6362
Author: rana 
Type: Update 
Category: SUS 
Subject: Optical Lever had bad filters 

 We found that that bounce (16.1 Hz) and roll (23.5 Hz) modes on the ITMX were much higher than on the ITMY. After some checking, it seems that the bandstop filters for the

SUSPOS, SUSPIT, SUSYAW, and SUSSIDE loops are set to the correct frequencies. However, the OLPIT and OLYAW had not been set correctly. I have copied the SUS filters into the OL filterbanks and reloaded all the filter banks. Attached are the comparison of old, bad, OL with the SUS ones.

The same cockamamie situation was there for the BS & ITMY as well. Although we still don't have the roll mode frequencies listed in the mechanical resonances wiki, I have guessed that the ITMY roll frequency is the same as the ITMX, since they have nearly the same bounce frequency. OL filters for the BS & ITMY are now at the right frequency (probably). Keiko is on top of fixing things for the other optics.

I think this whole notching adventure was in Leo's hands several months ago, but WE forgot to point him at the OLs in addition to the SUS. I blame Kiwamu 50% for not supervising him and Koji by 45% for not supervising Kiwamu. The other 5% goes to someone else. You know who you are.

Attachment 1: Untitled.png  448 kB  | Hide | Hide all
Untitled.png
ELOG V3.1.3-